OpenStudy (anonymous):

((4)/(5))^(x+5)=((25)/(16))^(-2x)

OpenStudy (anonymous):

(4/5)^(x+5)=(4/5)^4x 4x=x+5 x=5/3

OpenStudy (anonymous):

\[(4\div5)^{x+5} = (25\div16)^{-2x}\] \[(4\div5)^{x+5} = (16\div25)^{2x}\] \[(4\div5)^{x+5} = (4\div5)^{4x}\] \[x + 5 = 4x\]